# Scalable Multi-Task Learning for Particle Collision Event Reconstruction with Heterogeneous Graph Neural Networks

### v1

@@ 7 Conclusion
Reconstructing beauty hadron decays amidst the high-luminosity conditions of current and future LHC runs presents formidable challenges, including higher particle multiplicities and the prevalence of overlapping primary vertices (PVs), further compounded by strict latency and storage requirements. To address these issues, we have developed a novel HGNN architecture that jointly prunes background, reconstructs beauty-hadron decays, and associates tracks to PVs. By integrating pruning at every layer and optimising with a multi-objective loss, our HGNN achieves substantially higher beauty-hadron reconstruction efficiency and background rejection than the earlier DFEI multi-stage GNN framework [22], while limiting CPU/GPU inference time as particle multiplicity grows.
Central to our design are unique representations for particle tracks and candidate vertices, enabling the HGNN to learn their mutual relations and to assign each track to its true origin vertex with high accuracy. Early-layer pruning coupled with weighted message passing reduces the graph size with negligible performance loss, bounding inference latency even as particle multiplicity grows. These architectural innovations deliver substantial gains in PV association and beauty reconstruction performance, while moving us closer to the stringent latency and storage requirements at data acquisition.
The potential improvement in PV association and beauty hadron reconstruction, would translate into improved sensitivities of precision measurements across the LHCb physics program, reducing sources of background and improving resolutions of derived quantities. More broadly, the architectural novelties are applicable in several particle physics experiments, where heterogeneous datasets, multi-task reconstruction and the scalability of GNN inference time are common.
@@ 8 Acknowledgements
W. S. and N. S. have received support from the Swiss National Science Foundation (SNF) under TMAG-2_209263. J.G.P. is supported by the U.S. National Science Foundation under Grant No. 2411204. We acknowledge support from the Italian national funding agency INFN and CERN.
@@ Appendix A Training curves
The multi-task training optimises several objectives losses simultaneously. Figure 7 shows contributions to the overall training and validation loss in from equation 3 for GNN and HGNN trainings with training epochs.
@@ Appendix B Multi-task hyperparameter optimisation
The improvements in performance arising from the multi-task training is dependent on the $\beta$ parameters which govern the significance of the various tasks. To demonstrate this, we performed a two-dimensional grid scan in $\beta^{e_{\rm tr}}$ and $\beta^{v_{\rm tr}}$ as shown in Figure 8, which indicates that the relative significance of the edge and node pruning tasks influence the final $\mathcal{L}^{\rm LCA}_{\rm CE}$ value.
Refer to captionRefer to caption
Figure 8: Grid scan in $\beta^{e_{\rm tr}}$ and $\beta^{v_{\rm tr}}$ showing the final LCA reconstruction loss value, $\mathcal{L}^{\rm LCA}_{\rm CE}$. At the approximate position of the minima, $(\beta^{e_{\rm tr}},\beta^{v_{\rm tr}})=(40,0.6)$, a one-dimensional scan in $\beta^{e_{\rm pv}}$ is shown.
Each training is performed in a similar manner to the procedure described in Section 4.1. A global minimum is found approximately in the region $(\beta^{e_{\rm tr}},\beta^{v_{\rm tr}})=(40,0.6)$, which indicates that a relatively high weighting of the edge pruning task benefits the LCAG reconstruction task.
Appendix C Batch timing
Further accelerations in model inference time are possible by coupling batch inference with early layer pruning. Figure 9 shows the average inference time per event when processing batches of 8 events as a function of the average track multiplicity within the batch.
Figure 9: Mean GPU and CPU inference time per event with mean track multiplicity for batches of 8 events with and without early pruning ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$).
Appendix D PV association
Figure 10 compares the dependence of PV-association for beauty hadrons between the minimum impact parameter method and approaches using HGNNs.
Figure 10: Dependence of the average PV association (%) for beauty hadrons between various methods.

### v2

@@ 7 Conclusion
Reconstructing beauty hadron decays amid the high-luminosity conditions of current and future LHC runs presents formidable challenges, including higher particle multiplicities and the prevalence of overlapping primary vertices (PVs), further compounded by strict latency and storage requirements. To address these issues, we developed a novel HGNN architecture that jointly prunes background particles, reconstructs beauty-hadron decays, and associates tracks with their correct PV. By integrating pruning at every layer and optimising with a multi-objective loss, our HGNN achieves substantially higher beauty-hadron reconstruction efficiency and background rejection than the earlier DFEI multi-stage GNN framework [22], while limiting the CPU/GPU inference time as the particle multiplicity increases.
Central to our design are the unique representations for particle tracks and candidate vertices, enabling the HGNN to learn their mutual relations and assign each track to its true origin vertex with high accuracy. Early-layer pruning coupled with weighted message passing reduces the graph size with negligible performance loss, bounding inference latency even as particle multiplicity grows. These architectural innovations deliver substantial gains in PV association and beauty reconstruction performance, while moving closer to the stringent latency and storage requirements for data acquisition.
The potential improvement in PV association and beauty hadron reconstruction would translate into improved sensitivities of precision measurements across the LHCb physics program, reducing sources of background and improving resolutions of derived quantities. More broadly, the architectural novelties are applicable in several particle physics experiments, where heterogeneous datasets, multi-task reconstruction and the scalability of GNN inference time are common.
@@ 8 Acknowledgements
W. S. and N. S. received support from the Swiss National Science Foundation (SNF) under TMAG-2_209263. J.G.P. was supported by the U.S. National Science Foundation under Grant No. 2411204. Furthermore, we acknowledge support from the Italian national funding agency INFN and CERN.
9 Conflicts of Interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
10 Data availability statement
The data and code supporting the findings of this study are openly available at the following URLs: https://zenodo.org/records/15584745 and https://github.com/willsutcliffe/scalable mtl hgnn
@@ Appendix A Training curves
The multi-task training simultaneously optimises several objectives losses. Figure 7 shows the contributions to the overall training and validation loss in Equation 3 for the GNN and HGNN trainings as a function of training epochs.
@@ Appendix B Multi-task hyperparameter optimisation
The improvement in performance arising from multi-task training is dependent on the $\beta$ parameters, which govern the significance of the various tasks. To demonstrate this, we performed a two-dimensional grid scan in $\beta^{e_{\rm tr}}$ and $\beta^{v_{\rm tr}}$ for the GNN architecture as shown in Figure 8, which indicates that the relative scaling of the edge and node pruning tasks influences the final $\mathcal{L}^{\rm LCA}_{\rm CE}$ value.
Each training was performed in a manner similar to the procedure described in Section 4.1. A global minimum is found approximately in the region $(\beta^{e_{\rm tr}},\beta^{v_{\rm tr}})=(40,0.6)$, which indicates that a relatively high weighting of the edge pruning task benefits the LCAG reconstruction task.
Figure 8: Grid scan in $\beta^{e_{\rm tr}}$ and $\beta^{v_{\rm tr}}$ showing the final validation LCAG loss value, $\mathcal{L}^{\rm LCA}_{\rm CE}$.
Appendix C PV association
Figure 9 compares the dependence of the PV association for beauty hadrons between the minimum impact parameter method and the approaches using HGNNs.
Figure 9: Dependence of the average PV association (%) on PV multiplicity for beauty hadrons when using various HGNN variants and the minimum impact parameter method for PV association.